OpenStudy (anonymous):

A shopper purchased 8 t shirts and 5 pairs of pants for 220. The next day he purchased 5 t shirts and 1 pair of pants for 112. How much does each t shirt and each pair of pants cost?

OpenStudy (anonymous):

Set it up as a system of equations. Let x=t-shirts and y=pants

OpenStudy (anonymous):

eq1: \[8x+5y=220\] eq2: \[5x+y=112\] Eleminate one of the variables, which in this case "y" would be an easy choice. So, multiply eq.2 by -5 \[(5x+y=112)-5\] \[=-25x-5y=-560\] Now, add that to eq.1: \[8x+5y=220\] \[-25x-5y=-560\] = \[-17x=-340\] Solve for x, then substitute x into either original eq. and you'll get y